OpenStudy (anonymous):

What is the Value of h when the function is converted to vertex form? F(x)=x2+10x+35

OpenStudy (zpupster):

do you know what vertex form is??

OpenStudy (anonymous):

um i'm not sure

OpenStudy (zpupster):

this is the form \[=a(x-h)^2 +k\]

OpenStudy (zpupster):

(h,k) is the vertex

OpenStudy (zpupster):

the easiest way is : ax^2 + bx + c F(x)=x2+10x+35 h= -b/2a plug in b=10 a=1 so -10/2(1) what is h

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

h=-5

OpenStudy (zpupster):

yes that is your h value.